# Supernovae in 2023 (review): breakthroughs by late observations

### v3

@@ Abstract
I present a review of how late observations of supernovae, of the nebular phase, and much later of supernova remnants (SNRs), and their analysis in 2023 made progress towards breakthroughs in supporting the jittering jets explosion mechanism (JJEM) for core-collapse supernovae (CCSNe) and in introducing the group of lonely white dwarf (WD) scenarios for type Ia supernovae (SNe Ia). The new analyses of CCSN remnants (CCSNRs) reveal point-symmetric morphologies in a way unnoticed before in three CCSNRs. Comparison to multipolar planetary nebulae that are shaped by jets suggests that jets exploded these CCSNe, as predicted by the JJEM, but incompatible with the prediction of the delayed neutrino explosion mechanism. The spherical morphology of the ejecta Pa 30 of the historical type Iax supernova (SN Iax) of 1181 AD, which studies in 2023 revealed, is mostly compatible with the explosion of a lonely WD. Namely, at the explosion time, there is only a WD, without any close companion, although the WD was formed via a close binary interaction, i.e., binary merger. An identification of point-symmetry in SNR G1.9+0.3, a normal SN Ia and the youngest SN in the Galaxy, suggests an SN explosion of a lonely WD inside a planetary nebula (an SNIP). The group of lonely WD scenarios includes the core degenerate scenario and the double degenerate scenario with a merger to explosion delay (MED) time. SN Ia explosions of lonely WDs are common, and might actually account for most (or even all) normal SNe Ia.
@@ 1 Introduction
As such, this review focuses on papers from the year 2023 that are relevant to either the explosion mechanism of CCSNe or to the correspondence between different scenarios of SNe Ia and observed types of SNe Ia, including normal SNe Ia and peculiar SNe Ia. I will not review the basic physics and observations of these supernovae, nor will I review pair-instability SNe (e.g., Schulze et al. 2023). I will almost not cite papers before the year 2023 as these are cited in the relevant papers from 2023 that I cite. For extended reviews of CCSN explosion mechanisms, the reader should consult earlier reviews, e.g., Janka et al. (2016) and Mezzacappa (2023) for reviews of the physics of CCSNe and the delayed-neutrino explosion mechanism and Soker (2022) on the jittering jets explosion mechanism (JJEM). In Soker (2022) I argued that models that attribute the extra energy of super-luminous CCSNe in the frame of the delayed neutrino explosion mechanism to magnetars must include jets because jets during and/or after the explosion accompany the formation of energetic magnetars. I will return to this important point in Section 3.1.
The different scenarios of SNe Ia motivated many reviews on the topic of SNe Ia in the last decade which together cover all aspects of SNe Ia prior to 2023 (e.g., Maoz et al. 2014; Maeda, & Terada 2016; Hoeflich 2017; Livio & Mazzali 2018; Soker 2018, 2019; Wang 2018; Jha et al. 2019; Ruiz-Lapuente 2019; Ruiter 2020; Liu, Röpke, & Han 2023). Vinkó, Szalai, & Könyves-Tóth (2023) present a short review of some recent aspects of SNe Ia and CCSNe. For observational classification see, e.g., Aleo et al. (2023).
CCSNe and SNe Ia differ in many aspects from each other. They also have some common physical processes. The common ground of CCSNe and SNe Ia in this review is the usage of supernova remnants (SNRs). I will use the morphologies of some SNRs to reach what I consider to be significant steps towards breakthroughs in 2023. One breakthrough is in coming closer to answering the question of what is the explosion mechanism of CCSNe? (Section 3.2) and the other breakthrough is in establishing the group of lonely white dwarf (WD) scenarios of SNe Ia, which I argue hints at the dominant scenarios of normal SN Ia (Section 4.2). I start (Section 2) by describing the point-symmetric morphology that is relevant to the SNRs that I describe. I open each of the following two Sections, Section 3 on CCSNe and Section 4 on SNe Ia, with a review of relevant papers from 2023 and then present new results that I claim might be considered breakthroughs. I summarize this review in Section 5.
@@ 2 The point-symmetric morphology of SNRs
Because the study of SNR morphology is relevant both to CCSNe and to the progenitors of SNe Ia, I discuss the point-symmetry morphology here.
@@
The study of SNRs (their properties, their progenitors, their interaction with the ISM, particle acceleration, and more) is a subject by itself that continued to be a major topic in 2023 (e.g., Araya 2023; Bakış et al. 2023; Ball et al. 2023; Bamba et al. 2023; Banovetz et al. 2023; Bozzetto et al. 2023; Braun et al. 2023; Bykov et al. 2023; Caldwell & Raymond 2023; Cao et al. 2023; Chambery et al. 2023; Deng et al. 2023; Dokara et al. 2023; Domček et al. 2023; Duvidovich & Petriella 2023; Eagle et al. 2023; Ellien, Greco, & Vink 2023; Ferrazzoli et al. 2023; Fujishige et al. 2023; Ghavamian et al. 2023; Godinaud et al. 2023; Himono et al. 2023; Holland-Ashford, Slane, & Long 2023; Katsuda 2023; Khabibullin et al. 2023; Kim, Koo, & Onaka 2023; Koplitz et al. 2023; Makarenko et al. 2023; Meyer et al. 2023; Narita et al. 2023; Nguyen-Dang et al. 2023; Ranasinghe & Leahy 2023; Rho et al. 2023a, b; Sankrit, Blair, & Raymond 2023; Sano et al. 2023; Seward & Points 2023; Shanahan et al. 2023; Sharma et al. 2023a; Sinitsyna & Sinitsyna 2023; Suherli et al. 2023; Sushch & Brose 2023; Suzuki et al. 2023; Vasiliev & Shchekinov 2023; Winkler, Long, & Blair 2023; Winkler et al. 2023; Xiao et al. 2023; Xin & Guo 2023; Yamauchi & Pannuti 2023; Yeung, Bamba, & Sano 2023; Zhang et al. 2023b, c; Zhang & Xin 2023; Zhong et al. 2023; Zhou et al. 2023; Alsaberi et al. 2024a, b; Arias et al. 2024; Bao et al. 2024; Fesen et al. 2024; Khabibullin et al. 2024; Kirchschlager et al. 2024; López-Caraballo et al. 2024; Luo, Tang, & Mo 2024; Mantovanini et al. 2024; Michailidis et al. 2024a, b; Milisavljevic et al. 2024; Picquenot et al. 2024; Reyes-Iturbide et al. 2024; Reynolds & Borkowski 2024; Romano, Behrendt, & Burkert 2024; Sato et al. 2024; Vink et al. 2024; Zhang et al. 2024; Zhou et al. 2024). I will focus only on the morphologies of some SNRs.
@@
A point symmetric morphology is one where for each of the major morphological features in a nebula or a remnant there is an opposite morphological feature with respect to the center. Such morphological features might be clumps, filaments, lobes, and ears. An ear is a protrusion from the main nebula or remnant that has a base smaller than the main nebula/remnant and a cross-section that decreases outwards. Cotton et al. (2023) argue that ears are common in SNRs. This is also evident from the classification of CCSN remnants (CCSNRs; Soker 2023d).
@@
The point-symmetry morphology is the main SNR property that was used in 2023 to point to a possible breakthrough in our understanding of CCSN explosion mechanisms and to establish the group of lonely WD SN Ia scenarios. I make use of the large body of knowledge that exists in the study of bipolar and point-symmetric planetary nebulae, like multi-polar planetary nebulae. Many studies of planetary nebulae show that bipolar and point-symmetric structures are shaped by jets (e.g., Akashi 2023; Gómez-Muñoz et al. 2023; Lora et al. 2023; Mari, Gonçalves, & Akras 2023; Moraga Baez et al. 2023 for some 2023 studies). Comparison of some SNR morphologies to point-symmetric planetary nebulae can lead to the identification of point-symmetric SNRs (e.g., Soker 2023d, 2024c).
@@
The interaction of SN ejecta with a circumstellar material (CSM) or with the interstellar medium (ISM) cannot shape a point-symmetry morphology unless the ejecta and/or the CSM have point-symmetric morphology, to begin with; I here mention only studies from 2023 (see discussion in Soker 2024a). A magnetized medium cannot form a point-symmetry with three or more pairs of opposite ears/clumps/filaments/lobes as they form a structure with one symmetry axis, or rarely two symmetry axes, one of the exploding jets and one of the ambient magnetic field direction (e.g., Meyer et al. 2024). Velázquez et al. (2023) simulated a non-spherical wind interaction with a magnetized ISM. They find that in a pre-explosion axisymmetric wind with an axis inclined to the interstellar medium (ISM) magnetic field, the ears of the SNR might be bent. Namely, not exactly on a straight line through the center. However, this process cannot shape point-symmetric filaments/clumps/ears/lobes. Zhang et al. (2023d) included magnetic fields and ISM density gradients in their simulation of SNR G1.9+0.3 (that I study in Section 4.2.3). They could form an SNR with a pair of ears but not a point-symmetric SNR. Villagran et al. (2024) performed magneto-hydrodynamic simulations aiming at shaping SNR G1.9+0.3. They blew a non-spherical pre-explosion wind into a magnetized ISM. They obtained an axisymmetrical morphology, but not a point-symmetric morphology. Instabilities that are excited by ejecta-ISM interaction, as Mandal et al. (2023) study, are unable to form point-symmetric morphology. Turbulence in the interstellar medium (ISM), e.g., Rigon & Inoue (2023), also cannot form point-symmetric structures.
@@
My conclusion is that point-symmetric SNRs are shaped by jets. In this review, I present results from 2023 attributing the point-symmetric structures of three CCSNRs (N63A, Vela, SN 1987A) to point-symmetric ejecta shaped by jets that were part of the pairs of jets that exploded the star in the frame of the JJEM, while a point-symmetric SNR Ia (SNR G1.9+0.3) is attributed to a point-symmetric CSM which was a planetary nebula.
@@ 3.1 Neutrino-driven versus jet-driven explosions
I start by reviewing some (but not all) studies of CCSNe that are relevant to the quest for the explosion mechanism. In comparing theory with observations, one must bear in mind the large diversity of CCSNe, as studies in 2023 continued to show (e.g., Agudo et al. 2023; Ailawadhi et al. 2023; Aryan et al. 2023; Ben-Ami et al. 2023; Blanchard et al. 2023; Bostroem et al. 2023a, c; Chen et al. 2023b; Corsi et al. 2023; Desai et al. 2023a; Dong et al. 2023b; Dorn-Wallenstein, Neugent, & Levesque 2023; Gangopadhyay et al. 2023; Holmbo et al. 2023; Irani et al. 2023; Jeena, Banerjee, & Heger 2023; Jiang & Liu 2023; Jin, Yoon, & Blinnikov 2023; Kilpatrick et al. 2023b; Könyves-Tóth & Seli 2023; Kuncarayakti et al. 2023; Li, Zhong, & Dai 2023; Lin, Zhang, & Zhang 2023a; Matsuoka et al. 2023; Moore et al. 2023; Moriya et al. 2023c; Park, Yoon, & Blinnikov 2023; Pessi et al. 2023b; Pessi, Prieto, & Dessart 2023; Pursiainen et al. 2023b; Rodríguez, Maoz, & Nakar 2023; Scully, Matzner, & Yalinewich 2023; Shivkumar et al. 2023; Sit et al. 2023; Srinivasaragavan et al. 2023; Stritzinger et al. 2023a, b; Subrayan et al. 2023; Sun, Maund, & Crowther 2023; Teja et al. 2023; Utrobin & Chugai 2023; Van Dyk et al. 2023a; Vasylyev et al. 2023a; Vazquez et al. 2023; Wang et al. 2023a, b; Wen et al. 2023; Williamson et al. 2023; Yan et al. 2023; Zhu et al. 2023; Bjornsson 2023; Brennan et al. 2024b; Burrows, Wang, & Vartanyan 2024; Chugai & Utrobin 2024; Das et al. 2024; Dastidar et al. 2024; Dessart et al. 2024; Dessart, Hillier, & Leonard 2024b; Elias-Rosa et al. 2024; Kangas et al. 2024; Lin et al. 2024; Sato et al. 2024; Shahbandeh et al. 2024; Sharma et al. 2024).
@@
I refer here to two explosion mechanisms: the delayed neutrino explosion mechanism and the JJEM. There are many new studies in 2023 of different aspects of the delayed neutrino explosion mechanism (e.g., Aguilera-Dena et al. 2023; Barker, O’Connor, & Couch 2023; Boccioli et al. 2023; Burrows, Vartanyan, & Wang 2023; Gogilashvili, Murphy, & Miller 2023; Navó et al. 2023; Orlando 2023; Powell et al. 2023; Sieverding, Kresse, & Janka 2023; Sumiyoshi et al. 2023; van Baal et al. 2023; Wang & Burrows 2023a, b; Zha et al. 2023; Andresen et al. 2024; Ghodla & Eldridge 2023; Janka & Kresse 2024; Maunder et al. 2024; Schneider, Podsiadlowski, & Laplace 2024; Wang & Burrows 2024). Varma, Müller, & Schneider (2023), for example, find with their 3D simulations that magnetic fields help drive earlier and more energetic explosions (see also Matsumoto, Takiwaki, & Kotake 2024). There is a lower number of new studies of the JJEM (e.g., Antoni & Quataert 2023; Bear & Soker 2023; Shishkin & Soker 2023; Soker 2023e), including the operation of the JJEM in electron capture supernovae (Wang, Shishkin, & Soker, 2024), that potentially can take place for stars with an initial mass of $8.5M_{\odot}\lesssim M_{\rm ZAMS}\lesssim 9.2M_{\odot}$ (e.g., Guo et al. 2024; Limongi et al. 2023). I will devote Section 3.2 to describe new results that I consider to strongly support the JJEM, and are toward a breakthrough in answering the question of the CCSN explosion mechanism. I therefore hope this review will motivate more studies of the JJEM.
@@
Both of these explosion mechanisms require instabilities in the zone between the newly born neutron star (NS) and the stalled shock of the collapsing core at $r\simeq 100-150{~{}\rm km}$. The delayed neutrino explosion mechanism requires the instabilities to revive the stalled shock (e.g., Fryer et al. 2023). The JJEM requires the instabilities to increase the angular momentum seed perturbations that the convective zones of the collapsing core supply, in particular the spiral modes of the standing accretion shock instability (SASI). For a recent study of the spiral SASI see Buellet et al. (2023).

### v4

# Supernovae in 2023 (review): possible breakthroughs by late observations

@@ Abstract
I present a review of how late observations of supernovae, of the nebular phase, and much later of supernova remnants (SNRs), and their analysis in 2023 made progress towards possible breakthroughs in supporting the jittering jets explosion mechanism (JJEM) for core-collapse supernovae (CCSNe) and in introducing the group of lonely white dwarf (WD) scenarios for type Ia supernovae (SNe Ia). The new analyses of CCSN remnants (CCSNRs) reveal point-symmetric morphologies in a way unnoticed before in several CCSNRs. Qualitative comparison to multipolar planetary nebulae that are shaped by jets suggests that jets exploded these CCSNe, as predicted by the JJEM, but incompatible with the prediction of the delayed neutrino explosion mechanism. The spherical morphology of the ejecta Pa 30 of the historical type Iax supernova (SN Iax) of 1181 AD, which studies in 2023 revealed, is mostly compatible with the explosion of a lonely WD. Namely, at the explosion time, there is only a WD, without any close companion, although the WD was formed via a close binary interaction, i.e., binary merger. Identifying point-symmetry in SNR G1.9+0.3, a normal SN Ia and the youngest SN in the Galaxy, suggests an SN explosion of a lonely WD inside a planetary nebula (an SNIP). The group of lonely WD scenarios includes the core degenerate scenario and the double degenerate scenario with a merger to explosion delay (MED) time. SN Ia explosions of lonely WDs are common and might actually account for most (or even all) normal SNe Ia.
@@ 1 Introduction
As such, this review focuses on papers from the year 2023 that are relevant to either the explosion mechanism of CCSNe or to the correspondence between different scenarios of SNe Ia and observed types of SNe Ia, including normal SNe Ia and peculiar SNe Ia. I will not review the basic physics and observations of these supernovae, nor will I review pair-instability SNe (e.g., Schulze et al. 2023). I will almost not cite papers before the year 2023 as these are cited in the relevant papers from 2023 that I cite. For extended reviews of CCSN explosion mechanisms, the reader should consult earlier reviews, e.g., Janka et al. (2016) and Mezzacappa (2023) for reviews of the physics of CCSNe and the delayed-neutrino explosion mechanism and Soker (2022c) on the jittering jets explosion mechanism (JJEM). In Soker (2022c), I argued that models that attribute the extra energy of super-luminous CCSNe in the frame of the delayed neutrino explosion mechanism to magnetars must include jets because jets during and/or after the explosion accompany the formation of energetic magnetars. I will return to this important point in Section 3.1.
The different scenarios of SNe Ia motivated many reviews on the topic of SNe Ia in the last decade which together cover all aspects of SNe Ia before 2023 (e.g., Maoz et al. 2014; Maeda, & Terada 2016; Hoeflich 2017; Livio & Mazzali 2018; Soker 2018, 2019; Wang 2018; Jha et al. 2019; Ruiz-Lapuente 2019; Ruiter 2020; Liu, Röpke, & Han 2023). Vinkó, Szalai, & Könyves-Tóth (2023) present a short review of some recent aspects of SNe Ia and CCSNe. For observational classification, see, e.g., Aleo et al. (2023).
CCSNe and SNe Ia differ in many aspects from each other. They also have some common physical processes. The common ground of CCSNe and SNe Ia in this review is the usage of supernova remnants (SNRs). I will use the morphologies of some SNRs to reach what I consider to be significant steps towards breakthroughs in 2023. One breakthrough is coming closer to answering the question of what is the explosion mechanism of CCSNe? (Section 3.2) and the other breakthrough is in establishing the group of lonely white dwarf (WD) scenarios of SNe Ia, which I argue hints at the dominant scenarios of normal SN Ia (Section 4.2). I start (Section 2) by describing the point-symmetric morphology that is relevant to the SNRs that I describe. I open each of the following two Sections, Section 3 on CCSNe and Section 4 on SNe Ia, with a review of relevant papers from 2023 and then present new results that I claim might be considered breakthroughs. I summarize this review in Section 5.
@@ 2 The point-symmetric morphology of SNRs
Because the study of SNR morphology is relevant to CCSNe and to the progenitors of SNe Ia, I discuss the point-symmetry morphology here.
@@
The study of SNRs (their properties, their progenitors, their interaction with the ISM, particle acceleration, and more) is a subject by itself that continued to be a major topic in 2023 (e.g., Araya 2023; Bakış et al. 2023; Ball et al. 2023; Bamba et al. 2023; Banovetz et al. 2023; Bozzetto et al. 2023; Braun et al. 2023; Bykov et al. 2023; Caldwell & Raymond 2023; Cao et al. 2024; Chambery et al. 2023; Deng et al. 2023; Dokara et al. 2023; Domček et al. 2023; Duvidovich & Petriella 2023; Eagle et al. 2023; Ellien, Greco, & Vink 2023; Ferrazzoli et al. 2023; Fujishige et al. 2023; Ghavamian et al. 2024; Godinaud et al. 2023; Himono et al. 2023; Holland-Ashford, Slane, & Long 2024; Katsuda 2023; Khabibullin et al. 2023; Kim, Koo, & Onaka 2023; Koplitz et al. 2023; Makarenko et al. 2023; Meyer et al. 2023; Narita et al. 2023; Nguyen-Dang et al. 2023; Picquenot et al. 2023, 2024; Ranasinghe & Leahy 2023; Rho et al. 2023a, b; Sankrit, Blair, & Raymond 2023; Sano et al. 2023; Seward & Points 2023; Shanahan et al. 2023; Sharma et al. 2023a; Sinitsyna & Sinitsyna 2023; Suherli et al. 2024; Sushch & Brose 2023; Suzuki et al. 2023; Vasiliev & Shchekinov 2024; Winkler, Long, & Blair 2023; Winkler et al. 2023; Xiao et al. 2023; Xin & Guo 2023; Yamauchi & Pannuti 2023; Yeung, Bamba, & Sano 2023; Zhang et al. 2023b, c; Zhang & Xin 2023; Zhong et al. 2023; Zhou et al. 2023; Alsaberi et al. 2024a, b; Arias et al. 2024; Bao et al. 2024; Fesen et al. 2024; Khabibullin et al. 2024; Kirchschlager et al. 2024; López-Caraballo et al. 2024; Luo, Tang, & Mo 2024; Mantovanini et al. 2024; Michailidis et al. 2024a, b; Milisavljevic et al. 2024; Reyes-Iturbide et al. 2024; Reynolds & Borkowski 2024; Romano, Behrendt, & Burkert 2024; Sato et al. 2024; Vink et al. 2024; Zhang et al. 2024; Zhou et al. 2024). I will focus only on the morphologies of some SNRs.
@@
A point symmetric morphology is one where, for (almost) each of the major morphological features in a nebula or a remnant, there is an opposite morphological feature with respect to the center. Such morphological features might be clumps, filaments, lobes, and ears. An ear is a protrusion from the main nebula or remnant that has a base smaller than the main nebula/remnant and a cross-section that decreases outwards. Cotton et al. (2024) argue that ears are common in SNRs. This is also evident from the classification of CCSN remnants (CCSNRs; Soker 2023d).
@@
The point-symmetry morphology is the main SNR property that was used in 2023 to point to a possible breakthrough in our understanding of CCSN explosion mechanisms and to establish the group of lonely WD SN Ia scenarios. I use the large body of knowledge that exists in the study of bipolar and point-symmetric planetary nebulae, like multi-polar planetary nebulae. Many studies of planetary nebulae show that bipolar and point-symmetric structures are shaped by jets (e.g., Akashi 2023; Gómez-Muñoz et al. 2023; Lora et al. 2023; Mari, Gonçalves, & Akras 2023; Moraga Baez et al. 2023 for some 2023 studies). Comparison of some SNR morphologies to point-symmetric planetary nebulae can lead to the identification of point-symmetric SNRs (e.g., Soker 2023d, 2024c), as is the comparison to point-symmetric X-ray morphologies in groups and clusters of galaxies (Soker, 2024e).
@@
The interaction of SN ejecta with a circumstellar material (CSM) or with the interstellar medium (ISM) cannot shape a point-symmetry morphology unless the ejecta and/or the CSM have point-symmetric morphology, to begin with; I here mention only studies from 2023 (and early 2024; see discussion in Soker 2024a). A magnetized medium cannot form a point-symmetry with three or more pairs of opposite ears/clumps/filaments/lobes as they form a structure with one symmetry axis, or rarely two symmetry axes, one of the exploding jets and one of the ambient magnetic field direction (e.g., Meyer et al. 2024). Velázquez et al. (2023) simulated a non-spherical wind interaction with a magnetized ISM. They find that in a pre-explosion axisymmetric wind with an axis inclined to the interstellar medium (ISM) magnetic field, the ears of the SNR might be bent. Namely, not exactly on a straight line through the center. However, this process cannot shape point-symmetric filaments/clumps/ears/lobes. Zhang et al. (2023d) included magnetic fields and ISM density gradients in their simulation of SNR G1.9+0.3 (that I study in Section 4.2.3). They could form an SNR with a pair of ears but not a point-symmetric SNR. Villagran et al. (2024) performed magneto-hydrodynamic simulations aiming at shaping SNR G1.9+0.3. They blew a non-spherical pre-explosion wind into a magnetized ISM. They obtained an axisymmetrical morphology but not a point-symmetric morphology. Instabilities that are excited by ejecta-ISM interaction, as Mandal et al. (2023) study, cannot form point-symmetric morphology. Turbulence in the interstellar medium (ISM), e.g., Rigon & Inoue (2023), also cannot form point-symmetric structures.
@@
My conclusion is that point-symmetric SNRs are shaped by jets. In this review, I present results from 2023 (including early 2024) attributing the point-symmetric structures of three CCSNRs (N63A, Vela, SN 1987A; for five more point-symmetric CCSNRs see Soker 2024e) to point-symmetric ejecta shaped by jets that were part of the pairs of jets that exploded the star in the frame of the JJEM, while a point-symmetric SNR Ia (SNR G1.9+0.3) is attributed to a point-symmetric CSM which was a planetary nebula.
@@ 3.1 Neutrino-driven versus jet-driven explosions
I start by reviewing some (but not all) studies of CCSNe that are relevant to the quest for the explosion mechanism. In comparing theory with observations, one must bear in mind the large diversity of CCSNe, as studies in 2023 continued to show (e.g., Agudo et al. 2023; Ailawadhi et al. 2023; Aryan et al. 2023; Ben-Ami et al. 2023; Blanchard et al. 2023; Bostroem et al. 2023a, c; Chen et al. 2024a; Corsi et al. 2023; Das et al. 2023; Desai et al. 2023a; Dong et al. 2023b; Dorn-Wallenstein, Neugent, & Levesque 2023; Gangopadhyay et al. 2023; Holmbo et al. 2023; Irani et al. 2023; Jeena, Banerjee, & Heger 2024; Jiang & Liu 2023; Jin, Yoon, & Blinnikov 2023; Kilpatrick et al. 2023b; Könyves-Tóth & Seli 2023; Kuncarayakti et al. 2023; Li, Zhong, & Dai 2023; Lin, Zhang, & Zhang 2023a; Matsuoka et al. 2024; Moore et al. 2023; Moriya et al. 2023c; Park, Yoon, & Blinnikov 2024; Pessi et al. 2023b; Pessi, Prieto, & Dessart 2023; Pursiainen et al. 2023b; Rodríguez, Maoz, & Nakar 2023; Scully, Matzner, & Yalinewich 2023; Shivkumar et al. 2023; Sit et al. 2023; Srinivasaragavan et al. 2024; Stritzinger et al. 2023a, b; Subrayan et al. 2023; Sun, Maund, & Crowther 2023; Teja et al. 2023; Utrobin & Chugai 2024; Van Dyk et al. 2023a; Vasylyev et al. 2023a; Vazquez et al. 2023; Wang et al. 2023a, b; Wen et al. 2023; Williamson et al. 2023; Yan et al. 2023; Zhu et al. 2023; Bjornsson 2024; Brennan et al. 2024b; Burrows, Wang, & Vartanyan 2024; Chugai & Utrobin 2024; Das et al. 2024a; Dastidar et al. 2024; Dessart et al. 2024; Dessart, Hillier, & Leonard 2024b; Elias-Rosa et al. 2024; Kangas et al. 2024; Lin et al. 2024; Niu, Sun, & Liu 2024; Sato et al. 2024; Shahbandeh et al. 2024; Sharma et al. 2024).
@@
I refer here to two explosion mechanisms: the delayed neutrino explosion mechanism and the JJEM. There are many new studies in 2023 of different aspects of the delayed neutrino explosion mechanism (e.g., Aguilera-Dena et al. 2023; Barker, O’Connor, & Couch 2023; Boccioli et al. 2023; Burrows, Vartanyan, & Wang 2023; Gogilashvili, Murphy, & Miller 2024; Navó et al. 2023; Orlando 2023; Powell et al. 2023; Sieverding, Kresse, & Janka 2023; Sumiyoshi et al. 2023; van Baal et al. 2023; Wang & Burrows 2023a, b; Zha et al. 2023; Andresen et al. 2024; Ghodla & Eldridge 2023; Janka & Kresse 2024; Maunder et al. 2024; Schneider, Podsiadlowski, & Laplace 2024; Wang & Burrows 2024; for a recent review see Boccioli & Roberti 2024). Varma, Müller, & Schneider (2023), for example, find with their 3D simulations that magnetic fields help drive earlier and more energetic explosions (see also Matsumoto, Takiwaki, & Kotake 2024). There is a lower number of new studies of the JJEM (e.g., Antoni & Quataert 2023; Bear & Soker 2023; Shishkin & Soker 2023; Soker 2023e), including the operation of the JJEM in electron capture supernovae (Wang, Shishkin, & Soker, 2024), that potentially can take place for stars with an initial mass of $8.5M_{\odot}\lesssim M_{\rm ZAMS}\lesssim 9.2M_{\odot}$ (e.g., Guo et al. 2024; Limongi et al. 2024). I will devote Section 3.2 to describe new results that I consider to support the JJEM strongly and are toward a breakthrough in answering the question of the CCSN explosion mechanism. I therefore hope this review will motivate more studies of the JJEM.
A short description of the JJEM is in place here. The JJEM operates even if the pre-collapse core does not rotate. The idea is that instabilities above the newly born NS amplify stochastic angular momentum fluctuations that were formed in the convective layers of the pre-collapse core. The amplified angular momentum fluctuations are sufficiently large, according to the conjecture that needs future confirmation, to form stochastically varying intermittent accretion disks. These disks launch pairs of jets with varying directions and intensities. In total, during the explosion process that lasts for $\simeq 1-10{~{}\rm s}$ or even longer (e.g., Wang, Shishkin, & Soker 2024), there might be $\approx{\rm few}-30$ jet-launching episodes, each lasting for $\simeq 0.01-0.1\sec$ (Papish & Soker, 2014a). The jets launching velocities are $\simeq 10^{5}{~{}\rm km}{~{}\rm s}^{-1}$ (compatible with most CCSNe not having relativistic jets, e.g. Guetta et al. 2020). In each jet-launching episode, the jets carry a mass of $\approx 10^{-3}M_{\odot}$. These values imply that the newly born NS accretes a mass of $\approx 0.1M_{\odot}$ through intermittent accretion disks and that each accretion disk of an episode has a mass of $\approx 10^{-2}M_{\odot}$. Regarding the basic outcomes of the explosions, e.g., nucleosynthesis and lightcurves, for a given pre-explosion stellar model and the explosion energy, the JJEM’s expectation is that nucleosynthesis and lightcurves are very similar to those in the neutrino-driven mechanism. It is important to note that the JJEM does include heating by neutrinos but as a boosting process rather than the sole or the main effect (Soker, 2022b). In the case of energetic CCSNe beyond the reach of the models of the delayed-neutrino explosion mechanism, there is a need for dedicated simulations to calculate the outcomes since, in those cases, the JJEM requires more energetic jets. The major difference between the JJEM and the neutrino-driven explosion mechanism regarding the SN ejecta is the morphology of the ejecta, which is one of the major topics of this review and a new study (Soker 2024e).
The JJEM fundamentally differs (e.g., Soker 2022c) from the fixed-axis mechanism of the magnetorotational-driven CCSNe (e.g., Khokhlov et al. 1999; López-Cámara et al. 2013; Bromberg & Tchekhovskoy 2016; Gottlieb et al. 2022; Obergaulinger & Reichert 2023; Urrutia, De Colle, & López-Cámara 2023; Müller 2024): (1) According to the JJEM jets explode most, and possibly all, CCSNe, and not only rare CCSNe. (2) The JJEM operates even when the pre-collapse core does not rotate. (3) According to the JJEM there are no massive stars that end without explosion. Namely, there are no failed CCSNe. Even when a black hole is formed, there is an explosion driven by jets. (4) The JJEM operates in a jet-negative feedback mechanism: when the jets manage to expel the core in the explosion process, the accretion process of collapsing gas stops (with some delay time). This results in cases where the total explosion energy is several times the binding energy of the ejected mass.
@@
Both of these explosion mechanisms require instabilities in the zone between the newly born neutron star (NS) and the stalled shock of the collapsing core at $r\simeq 100-150{~{}\rm km}$. The delayed neutrino explosion mechanism requires the instabilities to revive the stalled shock (e.g., Fryer et al. 2023). The JJEM requires the instabilities to increase the angular momentum seed perturbations that the convective zones of the collapsing core supply, in particular, the spiral modes of the standing accretion shock instability (SASI). For a recent study of the spiral SASI, see Buellet et al. (2023).